OpenStudy (anonymous):

$8500, 1 year, 12%, monthly.

OpenStudy (lilg132):

8500 / 100 = $85 (this is 1 percent of 8500) to get 12% you must multiiply 85 by 12 85 * 12 = $1020 (This is 12% of 8500) now to work out how much they get in one year we multiply 1020 by 12 months 1020 * 12 = $12,240 per year dont forget to click good answer